### v1

# Kinetic formulation of compartmental epidemic models

Abstract
We introduce a kinetic model that couples the movement of a population of individuals with the dynamics of a pathogen in the same population. The model is formally shown to generalize the well-known compartmental models in mathematical epidemiology. We prove the existence and uniqueness of solutions in appropriate spaces for particular instances of the model. We finish with some examples, and discuss possible applications of this modeling approach.
Keywords: Kinetic models; macroscopic limits; compartimental models; existence and uniqueness of solutions.
MSC: 92D30, 92D25, 35R09.
1 Introduction
Kinetic models were introduced in the 19th century in the study of gas dynamics [Cer69, Cer88]; in the late 20th century, they found many applications in biological modeling, from which we highlight applications in chemotaxis [Pat53, ODA88, OS97, CMPS04, CDSM+06], immunology [DMMRRS22, ORRS24], cancer dynamics [BBD04, BD04], and, more recently, epidemiology [BDP21, BBDP21].
On the other hand, mathematical models of epidemiology have a long story that dates back to seminal works by Daniel Bernoulli in the 18th Century, followed closely by Jean d’Alembert. In these works, they studied the possible impact of variolation, a primitive technique akin to vaccination (but with some important differences); cf. [Bac11] for a detailed historical account with modern mathematical notation. After a long gap, in the early 20th Century, what is now known as compartmental models gained maturity. We refer to [Ros16, KMW27] for the original references, and [Thi03, Bri03, Mur02] for a modern approach.
In the present work, we consider the so-called SIRS model, in which the population is divided into three classes: the Susceptible, individuals that can be contaminated by a certain pathogen, the Infectious, individuals that can transmit the pathogen, and the Recovered, individuals that, after being infected, profit from a temporary or permanent immunity from the disease; when the immunity is permanent, it is customary to refer simply to the “SIR model”.
Transitions between the three classes above are given by $S+I\stackrel{{\scriptstyle\beta}}{{\to}}2I$, $I\stackrel{{\scriptstyle\gamma}}{{\to}}R$, and $R\stackrel{{\scriptstyle\alpha}}{{\to}}S$, where $\beta>0$, $\gamma>0$, and $\alpha\geq 0$, are the transmission rate, recovery rate, and loss of immunity rate, respectively.
One of the most well-known realizations of the above model is through the use of ordinary differential equations (ODE), in which case it reads
$\begin{cases}S^{\prime}&=\alpha R-\beta\frac{SI}{S+I+R}\ ,\\ I^{\prime}&=\beta\frac{SI}{S+I+R}-\gamma I\ ,\\ R^{\prime}&=\gamma I-\alpha R\ .\end{cases}$ (1)
This model is referred to as SIRS in the current work, including, by abuse of notation, the case $\alpha=0$. Several of its mathematical properties can be found in the references above.
We start by introducing a model inspired by the kinetic theory of gases, in which individuals move in a certain two-dimensional space; the movement is composed of a random reorientation, perturbed by a drift in a certain preferred direction that possibly depends on the full distribution of individuals (e.g., to avoid clusters of infectious individuals). Furthermore, depending on the encounters in state space (i.e, the space including not only position but also displacement velocity), individuals change their state from S to I, to R, and back to S, according to the SIRS dynamics discussed before. See Fig 1 for an illustration.
Figure 1: Illustration of the kinetic models for three classes of particles moving in a two-dimensional space. Each circle represents an individual, and the arrow indicates his or her velocity vector. Movements are composed of random reorientations perturbed by a density-dependent drift. Blue individuals are of S type, while red and green individuals are of I and R types, respectively. In the highlight, we see an encounter in phase space (i.e., same position in space and same velocity) of an S and an I individual, possibly resulting in two I individuals, according to the prescribed transition rate. Note that when two individuals have the same position and velocity, there is more opportunity (i.e., time) for the transmission to happen than if only the same position were required.
We benefit from two recently introduced ideas in the field of evolutionary games. In [AFMS21], a kinetic model is formulated with a transition kernel akin to the replicator dynamics in evolutionary dynamics, i.e., individuals change classes according to the value obtained, or payoff, in their interaction with the remaining individuals in the population. In this work, however, the velocity is not directly related to physical space movement, but rather an internal parameter. See [HS98] for a detailed description of evolutionary game dynamics, in particular of the replicator equation.
In [HC24], the SIRS evolutionary dynamics is written in the framework of evolutionary game theory, showing that the SIRS model can be reinterpreted as a mix of two different games, with two different group sizes.
Putting these two ideas together, we have a kinetic model with a transition kernel inspired by epidemic models.
There are several differences between the model introduced here and the one presented in [BDP21] that are worth a brief comment. The most important one is that we include class transitions directly inspired by the SIRS model, with the aim of having a macroscopic limit, when the problem is spatially homogeneous, the system (1). A second important difference is to consider more general movements of the individuals and proceed to a careful scale separation of the movement and the class transitions. Finally, we study the existence of solutions of the kinetic model, although in a limited manner. However, there are clear intersection points that could be explored in future works.
We finish this introduction with an outline of the present work. In Sec. 2 we introduce our basic model, i.e, the kinetic model for epidemiology; in Sec. 3 we formally show that its form macroscopic limit is, in the spatially homogeneous case, the SIRS model (1), validating the idea that the kinetic model is a generalization of the SIRS model. Furthermore, we define a mesoscopic model, a partial differential equation (PDE) of drift-diffusion-advection type, that lies in between the kinetic and the SIRS model. We proceed in Sec. 4 to show the existence of solutions in the appropriate spaces of particular instances of the kinetic and, by extension, of the PDE model derived before. We show three examples in Sec. 5 and discuss possible limitations and continuations of the present work in Sec. 6.
2 The Model
Let $f_{j}:\mathbb{R}_{+}\times\Omega\times V$, $j=S,I,R$, be the density of individuals of class $j$ at time $t\geq 0$, position $x\in\Omega\subset\mathbb{R}^{2}$ and velocity $v\in V\subset\mathbb{R}^{2}$. We assume that the space domain $\Omega$ is a bounded open set with smooth boundary and $V$ is compact and spherically symmetrical (in particular, $\int_{V}v\mathrm{d}v=0$). We define $|V|:=\int_{V}\mathrm{d}v$. Note that both $\Omega$ and $V$ are class-independent (i.e., do not depend on $j$); however, check Subsec. 5.3.
The density $f_{j}$ evolves according to a Boltzmann-type equation:
$\partial_{t}f_{j}+v\cdot\nabla_{x}f_{j}=\mathfrak{Q}_{j}[\mathbf{f}],\quad\mathbf{f}:=(f_{S},f_{I},f_{R})\ ,\quad j=S,I,R\ .$ (2)
Boundary conditions are given by $\hat{n}\cdot\nabla f_{j}\Big{|}_{\partial\Omega}=0$, $j=S,I,R$, where $\hat{n}$ is the unit exterior normal to the domain $\Omega$.
In the sequel, whenever we refer to results valid to all classes of individuals, we use the subindex $j$, without further specifications. We use the notation $f_{j}:=f_{j}(t,x,v)$, and $f^{\prime}_{j}:=f_{j}(t,x,v^{\prime})$.
The transition kernel $\mathfrak{Q}_{j}$ describes three different behaviors for the individual in the population: i) a random movement, associated to a mean regression with respect to the equilibrium distribution; ii) a preferred direction in which individuals tends to move towards more favorable environments; and iii) a class-transition kernel accounting to the pathogen spread dynamics, healing and immunity waning, according to the SIRS model.
More specifically, we assume
$\mathfrak{Q}_{j}[\mathbf{f}]=\mathcal{Q}_{j}^{(R)}[\mathbf{f}]+\mathcal{Q}_{j}^{(P)}[\mathbf{f}]+\mathcal{Q}^{(T)}_{j}[\mathbf{f}]\ ,$
where
  • The random reorientation kernel is given by
    $\mathcal{Q}_{j}^{(R)}[\mathbf{f}]=\lambda_{j}(t,x)\left(\frac{1}{|V|}\int_{V}f_{j}(t,x,v^{\prime})\mathrm{d}v^{\prime}-f_{j}(t,x,v)\right)\ ,\quad j=S,I,R\ .$
    The reorientation rate is given by $\lambda_{j}>0$. This particular expression means that, given a point $x\in\Omega$, the number of individuals with direction $v$ tends to increase if it is below average and decrease otherwise. We note that the kernel $\mathcal{Q}_{j}^{(R)}$ is conservative, i.e., $\int_{V}\mathcal{Q}_{j}^{(R)}[\mathbf{f}]\mathrm{d}v=0$. This simple fact will be useful in the subsequent calculations.
  • The preferred direction kernel models the rate of change to and from preferable and undesirable directions, respectively. We assume that only susceptible individuals have preferred directions and define
    $\mathcal{Q}_{S}^{(P)}[f_{I},f_{S}]$ $:=\mathcal{Q}_{S}^{(P)}[\mathbf{f}]$
    $=\int_{V}\bigl{(}T[f_{I}](v,v^{\prime})f_{S}(t,x,v^{\prime})-T[f_{I}](v^{\prime},v)f_{S}(t,x,v)\bigr{)}\mathrm{d}v^{\prime},$
    where the transition from direction $v^{\prime}$ to $v$ is given by
    $T[f_{I}](v,v^{\prime})=-\eta(t,x)v\cdot\nabla_{x}f_{I}+\xi(t,x)v^{\prime}\cdot\nabla_{x}f_{I}^{\prime}\ ,$
    with $\eta,\xi\geq 0$. In particular, this means that susceptible individuals try to avoid the direction with increasing concentrations of I individuals (first term in the right-hand side) and to keep the direction if it points toward a decrease of the concentration of I individuals. For completeness, we also define preferred direction kernels for I and R classes to be identically null:
    $\mathcal{Q}_{I}^{(P)}[\mathbf{f}]:=\mathcal{Q}_{R}^{(P)}[\mathbf{f}]:=0\ ,\quad\forall\enspace\mathbf{f}\ .$
    It is straightforward to prove that the kernel $\mathcal{Q}^{(P)}_{j}$ is also conservative.
  • Finally, we define the class transition kernel, where
    $\mathcal{Q}_{j}^{(T)}[\mathbf{f}]=\Psi_{j}(\mathbf{f})f_{j}(t,x,v)\ ,\quad j=S,I,R\ ,$
    with
    $\Psi_{S}(\mathbf{f})$ $=-\beta\frac{f_{I}}{f_{S}+f_{I}+f_{R}}+\alpha\frac{f_{R}}{f_{S}}\ ,$
    $\Psi_{I}(\mathbf{f})$ $=\beta\frac{f_{S}}{f_{S}+f_{I}+f_{R}}-\gamma\ ,$
    $\Psi_{R}(\mathbf{f})$ $=\gamma\frac{f_{I}}{f_{R}}-\alpha\ .$
    Parameters $\alpha,\beta,\gamma$ may depend on $t$, $x$, and $v$, except if otherwise stated. We assume that transmission happens when one S and one I individual are aligned in space-phase (i.e., same position and same velocity), which gives more opportunity (i.e., more time) for the pathogen transmission. This is the reason we assume that the class-transition kernel depends on the phase-space density $f_{j}$ and not on the spatial densities $\int_{V}f_{j}\mathrm{d}v$.
The conservative property of the kernels $Q^{(R)}_{j}$ and $Q^{(P)}_{j}$ is not true in general, as class transitions are a fundamental feature of the proposed model, and the kernels $Q^{(T)}_{j}$ do not satisfy it. However, as we will see, the non-conservative part of the kernel is of higher order, and it will not introduce extra difficulties in the calculations of the formal limits.
Remark 1.
Assuming that $\Psi_{j}$ represents the relative fitness of type $j$ in a given population, it is clear that the average fitness is zero $\overline{\Psi}=f_{S}\Psi_{S}+f_{I}\Psi_{I}+f_{R}\Psi_{R}=0$. The choice of the expression is such that the replicator system of differential equations associated to the fitnesses functions $\Psi_{S}$, $\Psi_{I}$, and $\Psi_{R}$ is given by $f_{j}^{\prime}=f_{j}\left(\Psi_{j}-\overline{\Psi}\right)$, $\overline{\Psi}=\sum_{j}f_{j}\Psi_{j}$, is the SIRS system of differential equations. For further details, see [HC24].
See Fig. 2 to an illustration with the numeric implementation of the model.
Figure 2: Representation of the mesoscopic density $\rho_{S}$ (blue, left), $\rho_{I}$ (red, middle), and $\rho_{R}$ (right, green). More intense colors indicate higher concentrations. We consider the initial condition given by S and I individuals in the center of a square lattice of size $50\times 50$, with random initial velocities. Initial conditions are far from the boundary, and for times considered, they were irrelevant. Initially, there are 10000 S individuals and 50 I individuals. We plot the configuration after 3, 6, 9, 12, and 15 steps, each step including the update of all particles. Class-transition parameters were given by $\alpha=0$ (i.e., immunity is permanent), $\beta=0.75$ (each encounter between S and I individuals has a probability of $\nicefrac{{3}}{{4}}$ to generate a new I individual), and $\gamma=0.5$ (i.e., in each time step, each I individual has a probability $\nicefrac{{1}}{{2}}$ to became R. The random reorientation parameter is $\lambda=0.5$ and we did not consider preferred directions, i.e., $Q^{(P)}_{j}\equiv 0$.
3 The macroscopic limit
In this section, we study the macroscopic limit of the kinetic model in three steps. In Subsec. 3.1 we rewrite Eq. (2) in non-dimensional form, and identify small parameters; in the sequel, Subsec. 3.2, we identify an advection-reaction-diffusion partial differential equation (PDE) for the evolution of the density of individuals in the domain $\Omega$. We conclude in Subsec. 3.3, that spatially homogeneous solutions of the previous system of PDEs, whenever they exist, correspond to the solution of the SIRS epidemic model (1).
3.1 Adimensionalization of the kinetic model
In order to provide the dimensionless form of Eq. (2), we define typical observation times, spaces, velocities, and densities $t_{0}$, $x_{0}$, $v_{0}$, $f_{0j}$, respectively. We define new adimensional variables $\tilde{t}:=t/t_{0}$, $\tilde{x}:=x/x_{0}$, $\tilde{v}:=v/v_{0}$, and $\tilde{f}_{j}:=f_{j}/f_{0j}$ such that
$\partial_{\tilde{t}}\tilde{f}_{j}+\frac{t_{0}v_{0}}{x_{0}}\tilde{v}\cdot\nabla_{\tilde{x}}\tilde{f}_{j}=\frac{t_{0}}{f_{0j}}\mathfrak{Q}_{j}[\mathbf{f}_{0}\tilde{\mathbf{f}}]$
We note that $\frac{t_{0}v_{0}}{x_{0}}$ and $\frac{t_{0}}{f_{j0}}\mathfrak{Q}_{j}$ are non-dimensional. Furthermore, the microscopic velocity $v_{0}$ is much larger than the macroscopic velocity $x_{0}/t_{0}$, i.e.,
$\varepsilon:=\left(\frac{v_{0}}{x_{0}/t_{0}}\right)^{-1}\ll 1\ .$
We assume that typical density is low (an assumption akin to the rarefied gas assumption used in taking macroscopic limits of ideal gases modeled by Boltzmann equations). More precisely, we redefine the transition kernels $\varepsilon^{-2}\tilde{\mathfrak{Q}}_{j}[\mathbf{f}]=\frac{t_{0}}{f_{0j}}\mathfrak{Q}_{j}[\mathbf{f}_{0}\tilde{\mathbf{f}}]$, and omit the $\tilde{}$ notation to find the non-dimensional form of the kinetic equation
$\varepsilon^{2}\partial_{t}f_{j,\varepsilon}+\varepsilon v\cdot\nabla_{x}f_{j,\varepsilon}=\mathfrak{Q}_{j,\varepsilon}[\mathbf{f}_{\varepsilon}],$ (3)
where $\varepsilon$ is a small positive dimensionless parameter.
After proper adimensionalization, the transition kernel will also depend on the parameter $\varepsilon$. As we will see in this section, this implies in particular, that the three different effects discussed in Sec. 3 must differ in scales of $\varepsilon$ in order to find the correct macroscopic limit. More precisely, we assume the formal asymptotic of the kernel
$\mathfrak{Q}_{j,\varepsilon}=\mathcal{Q}_{j}^{(R)}+\varepsilon\mathcal{Q}_{j}^{(P)}+\varepsilon^{2}\mathcal{Q}_{j}^{(T)}\ .$
This means, in particular, that the leading order effect in the individual movement is the reorientation, followed by a small-order perturbation that gives preferred directions (when it is non-zero), and a second-order effect guaranteeing the class transitions are compatible with the SIRS model.
In the next subsection, we start to derive a partial differential equation for the purely spacial dependence of the leading order term, i.e., to
$\rho_{j}:=\lim_{\varepsilon\to 0}\int_{V}f_{j,\varepsilon}\mathrm{d}v\ .$
3.2 The mesoscopic limit
Before proceeding to obtain the mesoscopic limit, we remind the reader of a well-known theorem, whose proof can be found in the references:
Theorem 1 (Lax-Milgram Theorem [Ram80, Sho97, BCST25]).
[…]
In the sequel, we will use a more specific form of the Lax-Milgram theorem [^3] There is a slight confusion in the notation here, as the definition of the reorientation kernel was done in boldface variables. Here, we also omit the subindex $j$. the kernel $Q^{(R)}_{j}[\mathbf{f}]$, as defined in Sec. 2, depends only on $f_{j}$, and for the proof of the corollary we write only $Q^{(R)}[f]$ for $Q^{(R)}_{j}[\mathbf{f}]$. We hope this causes no confusion.:
Corollary 1.
Let $\mathcal{H}_{0}=\{f\in L^{2}(V)|\int_{V}f\mathrm{d}v=0\}$. Then, for each $g\in\mathcal{H}_{0}$ there is a unique $f\in\mathcal{H}_{0}$ such that $Q^{(R)}[f]=-g$, in the weak sense.
Proof.
It is clear that $\mathcal{H}_{0}$ is a Hilbert space with the natural $L^{2}$-inner product, $\langle f,g\rangle:=\int_{V}fg\mathrm{d}v$. Define the bilinear form $\mathcal{B}(g,h):=-\int_{V}Q^{(R)}[g]h\mathrm{d}v$, and
$\mathcal{B}(g,g)=-\lambda_{j}\left(\int_{V\times V}\frac{1}{|V|}g^{\prime}g\mathrm{d}v^{\prime}\mathrm{d}v-\int_{V}g^{2}\mathrm{d}v\right)=\lambda_{j}\langle g,g\rangle.$
From the fact that $\lambda_{j}$ does not depend on $v$ and is strictly positive, we conclude that $\mathcal{B}$ is coercive. The continuity is immediate.
The Lax-Miligram Theorem 1 guarantees that for each $g\in\mathcal{H}_{0}$, there is a unique $f\in\mathcal{H}_{0}$ such that
$\langle-g,h\rangle=-\mathcal{B}(f,h)=\int_{V}Q^{(R)}[f]h\mathrm{d}v\ ,\quad\forall h\in\mathcal{H}_{0}\ ,$
which finishes the proof. ∎
We assume the formal expansion of the solution
$f_{j,\varepsilon}=f_{j,0}+\varepsilon f_{j,1}+\varepsilon^{2}f_{j,2}\ .$
Note that $f_{j,1}$, $f_{j,2}$ are not particular instances of $f_{j,\varepsilon}$.
The objective of this subsection is to find a system of PDEs for $\rho_{j}=\int_{V}f_{j,0}\mathrm{d}v$.
Remark 2.
Although the notation can be confusing at this point, we call the reader’s attention to the fact that $f_{j,0}$ is the leading order expansion of $f_{j}$ and not the density normalization $f_{0j}$ introduced before.
Expanding both sides of Eq. (3) we find
$\varepsilon^{2}\partial_{t}f_{j,0}+\varepsilon v\cdot\nabla_{x}f_{j,0}+\varepsilon^{2}v\cdot\nabla_{x}f_{j,1}$
$\quad=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{0}]+\varepsilon\left(\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(P)}[\mathbf{f}_{0}]\right)$
$\qquad+\varepsilon^{2}\left(\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{2}]+\mathcal{Q}_{j}^{(P)}[\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(T)}[\mathbf{f}_{0}]\right)\ .$
In the remainder of this subsection, we compare terms of the same order of magnitude on both sides of the previous equation.
3.2.1 Leading Order Equation
Equating terms of order $\varepsilon^{0}=1$, we find
$0=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{0}]=\lambda_{j}\left(\frac{1}{|V|}\int_{V}f_{j,0}^{\prime}\mathrm{d}v^{\prime}-f_{j,0}\right)\ .$
The solution of the previous equation is such that
$f_{j,0}=\frac{1}{|V|}\int_{V}f_{j,0}^{\prime}\mathrm{d}v^{\prime}\ .$
We conclude that for all $j$, $\int_{V}vf_{j,0}\mathrm{d}v=0$, and this implies, in particular, that $f_{j,0}$ is spherically symmetric, i.e., there is a function $F:V\to\mathbb{R}_{+}$, with $F(v)=F(|v|)$, such that
$f_{j,0}(t,x,v)=\rho_{j}(x,t)F(|v|)\ ,$ (4)
and $\int_{V}F(v)\mathrm{d}v=1$.
3.2.2 Order 1 Equation
Equating terms of order $\varepsilon^{1}=\varepsilon$, we find
$v\cdot\nabla_{x}f_{j,0}=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(P)}[\mathbf{f}_{0}]$
For $j=I,R$, we only need to solve the equation
$\lambda_{j}\left(\frac{1}{|V|}\int_{V}f_{j,1}^{\prime}\mathrm{d}v^{\prime}-f_{j,1}\right)=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{1}]=Fv\cdot\nabla_{x}\rho_{j}\ .$ (5)
The solution is given by
$f_{j,1}=-\kappa_{j}(t,x,v)\nabla_{x}\rho_{j}(x,t)+\chi_{j}(t,x)F(v)\ .$ (6)
More precisely, from Corolary 1, it follows that there is $\bm{\kappa}(t,x,v)$ solution of
$\mathcal{Q}_{j}^{(R)}[\bm{\kappa}]=-vF\ ,\quad\bm{\kappa}=(\kappa_{S},\kappa_{I},\kappa_{R})\ ,\quad j=S,I,R\ ,$ (7)
and $\chi_{j}F$ is in the kernel of the operator $\mathcal{Q}_{j}^{(R)}$.
For $j=S$, we need to add an extra term on the right-hand side of the Eq. (5):
$Q_{S}^{(P)}[\mathbf{f}_{0}]$ $=\int_{V}\Big{(}-\eta(t,x)v\cdot(\nabla_{x}f_{I,0})f_{S,0}^{\prime}+\xi(t,x)v^{\prime}\cdot(\nabla_{x}f_{I,0}^{\prime})f_{S,0}^{\prime}$
$\quad+\eta(t,x)v^{\prime}\cdot(\nabla_{x}f_{I,0}^{\prime})f_{S,0}-\xi(t,x)v\cdot(\nabla_{x}f_{I,0})f_{S,0}\Big{)}\mathrm{d}v^{\prime}$
$=\int_{V}\Big{(}-\eta FF^{\prime}v\cdot(\nabla_{x}\rho_{I})\rho_{S}+\xi F^{\prime 2}v^{\prime}\cdot(\nabla_{x}\rho_{I})\rho_{S}$
$\quad+\eta FF^{\prime}v^{\prime}\cdot(\nabla_{x}\rho_{I})\rho_{S}-\xi F^{2}v\cdot(\nabla_{x}\rho_{I})\rho_{S}\Big{)}\mathrm{d}v^{\prime}$
$=-(\eta+\xi F|V|)vF\cdot(\nabla_{x}\rho_{I})\rho_{S}\ ,$
where $\eta=\eta(t,x)$, $\xi=\xi(t,x)$, and $F^{\prime}=F(v^{\prime})=F(|v^{\prime}|)$. Note that one particular realization of the leading order equilibrium distribution $F$ is $F(v)=\frac{1}{|V|}$. This will be the case whenever $V$ is a hypersphere (a circle, in the two-dimensional case). In this case, the last equation simplifies to $Q_{S}^{(P)}[\mathbf{f}_{0}]=-(\eta+\xi)vF\cdot(\nabla_{x}\rho_{I})\rho_{S}$.
We conclude that
$f_{S,1}=-\Theta(t,x,v)(\nabla_{x}\rho_{I})\rho_{S}-\kappa_{S}(t,x,v)\nabla_{x}\rho_{S}+\chi_{S}F$ (8)
where $\Theta(t,x,v)$ is the solution of $\mathcal{Q}_{S}^{(R)}[\Theta]=-(\eta+\xi)vF$, cf. Corollary 1.
3.2.3 Order 2 equations
The order $\varepsilon^{2}$ equation is given by
$\partial_{t}f_{j,0}+v\cdot\nabla_{x}f_{j,1}=\mathcal{Q}_{j}^{(R)}[\mathbf{f}_{2}]+\mathcal{Q}_{j}^{(P)}[\mathbf{f}_{1}]+\mathcal{Q}_{j}^{(T)}[\mathbf{f}_{0}]\ .$
Integrating on $V$ and using Eqs. (4) using the conservative property of $Q^{(P)}_{j}$, we find
$\partial_{t}\rho_{j,0}+\int_{V}v\cdot\nabla_{x}f_{j,1}\mathrm{d}v=\int_{V}\mathcal{Q}_{j}^{(T)}[\mathbf{f_{0}}]\mathrm{d}v\ .$
For $j=I$, we find
$\partial_{t}\rho_{I}-\int_{V}v\cdot\nabla_{x}\left[\kappa_{I}(t,x,v)\nabla_{x}\rho_{I}\right]\mathrm{d}v+\int_{V}v\cdot\nabla_{x}\left[\chi_{I}F(v)\right]\mathrm{d}v$
$\qquad=\int_{V}\left(\beta\frac{\rho_{S}}{\rho_{S}+\rho_{I}+\rho_{R}}-\gamma\right)\rho_{I}F(v)\mathrm{d}v=\beta_{0}\frac{\rho_{S}\rho_{I}}{\rho_{S}+\rho_{I}+\rho_{R}}-\gamma_{0}\rho_{I}\ ,$
where $\beta_{0}:=\int_{V}\beta F(v)\mathrm{d}v$, $\gamma_{0}:=\int_{V}\gamma F(v)\mathrm{d}v$; we define a similar equation for the loss of immunity parameter: $\alpha_{0}:=\int_{V}\alpha F(v)\mathrm{d}v$. Due to the normalization $\int_{V}F(v)\mathrm{d}v=1$, this means that, for each $x$ and $t$, $\alpha_{0}$, $\beta_{0}$, and $\gamma_{0}$ are averages at equilibrium of the corresponding parameters in the $V$-space.
Note that $\int_{V}v\cdot\nabla_{x}\left[\rho_{I}F(v)\right]\mathrm{d}v=\int_{V}vF(v)\mathrm{d}v\cdot\nabla_{x}\rho_{I}=0$.
Defining $D(t,x):=\int_{V}v\cdot\kappa(t,x,v)\mathrm{d}v$, we conclude
$\partial_{t}\rho_{I}=\nabla_{x}\left[D(t,x)\nabla_{x}\rho_{I}\right]+\beta_{0}\frac{\rho_{S}\rho_{I}}{\rho_{S}+\rho_{I}+\rho_{R}}-\gamma_{0}\rho_{I}\ .$ (9)
The equation for $j=R$ is similar:
$\partial_{t}\rho_{R}=\nabla_{x}\left[D(t,x)\nabla_{x}\rho_{R}\right]+\int_{V}\mathcal{Q}_{j}^{(T)}[\mathbf{f}_{0}]\mathrm{d}v=\nabla_{x}\left[D(t,x)\nabla_{x}\rho_{R}\right]+\gamma_{0}\rho_{I}-\alpha_{0}\rho_{R}\ .$ (10)
Finally, we study the equation for $j=S$. The only difference from the previous case is the existence of an extra term in Eq. (8) when compared to Eq. (6):
$\int_{V}v\cdot\nabla_{x}\left[\Theta(t,x,v)(\nabla_{x}\rho_{I})\rho_{S}\right]\mathrm{d}v=\nabla_{x}\cdot\left[\Gamma(t,x)(\nabla_{x}\rho_{I})\rho_{S}\right]\ ,$
where $\Gamma(t,x)=\int_{V}v\cdot\Theta(t,x,v)\mathrm{d}v$. Finally,
$\partial_{t}\rho_{S}=\nabla_{x}\left[D(t,x)\nabla_{x}\rho_{S}+\Gamma(t,x)\left(\nabla_{x}\rho_{I}\right)\rho_{S}\right]+\alpha_{0}\rho_{I}-\beta_{0}\frac{\rho_{S}\rho_{I}}{\rho_{S}+\rho_{I}+\rho_{R}}\ .$ (11)
Remark 3.
Assume that the transmission and recovery rates are such that an outbreak is locally prevented for any point in $\Omega$, for any time $t$, i.e., $\beta_{0}<\gamma_{0}$ for $(t,x)\in\mathbb{R}_{+}\times\Omega$. Multiplying Eq. (9) by $\rho_{I}$ and integrating in $\Omega$, we find
$\frac{1}{2}\frac{\mathrm{d}\ }{\mathrm{d}t}\int_{\Omega}\rho_{I}^{2}\mathrm{d}x\leq-\int_{\Omega}D(t,x)|\nabla_{x}\rho_{I}|^{2}\mathrm{d}x-\int_{\Omega}(\gamma_{0}-\beta_{0})\rho_{I}^{2}\mathrm{d}x<0\ ,$
where we used that $\frac{\rho_{S}}{\rho_{S}+\rho_{I}+\rho_{R}}\leq 1$. We conclude that $\rho_{I}\to 0$ for all $x\in\Omega$ when $t\to\infty$, i.e., the system approaches asymptotically the disease-free state. The condition on $\beta_{0}$ and $\gamma_{0}$ could be weakened without losing the disease-free state’s uniqueness as the only globally asymptotic attractor of the dynamics.
3.3 The SIRS epidemic model
Let us assume a space homogeneous problem, i.e., we consider that $\alpha_{0}$, $\beta_{0}$, and $\gamma_{0}$ do not depend on $x$, and that the initial condition is also space independent. Therefore, there is a space homogeneous solutions of the system (9)–(11), such that $(\rho_{S},\rho_{I},\rho_{R})$ is the solution of SIRS model 1, with $S=\rho_{S}$, $I=\rho_{I}$, and $R=\rho_{R}$.
However, when working with the SIRS model, we are interested in the total number of type S, I, and R individuals (possibly normalized to the total population), and not its average space-density. Therefore, it is natural to define
$S=\int_{\Omega}\rho_{S}\mathrm{d}x\ ,\quad I=\int_{\Omega}\rho_{I}\mathrm{d}x\ ,\quad R=\int_{\Omega}\rho_{R}\mathrm{d}x\ .$
$(S,I,R)$ is a solution of the SIRS model for constant-in-space $\rho_{j}$ with parameters
$\bar{\alpha}$ $:=\frac{1}{|\Omega|}\int_{\Omega}\alpha_{0}\mathrm{d}x=\frac{1}{|\Omega|}\int_{\Omega\times V}\alpha F(v)\mathrm{d}v\mathrm{d}x=\alpha_{0}\ ,$
$\bar{\beta}$ $:=\frac{1}{|\Omega|}\int_{\Omega}\beta_{0}\mathrm{d}x=\frac{1}{|\Omega|}\int_{\Omega\times V}\beta F(v)\mathrm{d}v\mathrm{d}x=\beta_{0}\ ,$
$\bar{\gamma}$ $:=\frac{1}{|\Omega|}\int_{\Omega}\gamma_{0}\mathrm{d}x=\frac{1}{|\Omega|}\int_{\Omega\times V}\gamma F(v)\mathrm{d}v\mathrm{d}x=\gamma_{0}\ .$
We finish this (small) subsection, noting that it is not difficult to generalize all the above processes for more general class transition kernels. In particular, it is possible to choose class transition kernels
$Q_{j}^{(T)}[\mathbf{f}]=f_{j}\left(\Psi_{j}[\mathbf{f}]-\sum_{j}f_{j}\Psi_{j}[\mathbf{f}]\right)\ ,$
where $\Psi_{j}[\mathbf{f}]$ is the fitness (reproductive viability) function for a population at state $\mathbf{f}$ (i.e, the presence of the type $j$ individuals is given by $f_{j}$. In this case, $\dot{f}_{j}=Q_{j}^{(T)}$ is the well-known replicator equation [HS98].
From the discussion in this section, the kinetic formulation is one possible generalization of the replicator equation, including not only space-inhomogeneity but also other variables (in particular, in this work, velocity).
4 Existence and uniqueness of solutions of the kinetic model
In this section, we prove the existence of solutions of the kinetic model (3), for fixed $\varepsilon>0$ (that will be omitted along this section), under the assumption that there is no preferred direction of movement, i.e., $\eta=\xi=0$. This assumption follows from the technical difficulty to prove existence and uniqueness results in kinetic modeling when transition kernels depend on the gradient of solutions, cf. [CMPS04, Remark 4]; see also [HS00] for the one-dimensional case with dependence on the gradient.
Theorem 2.
Let $\alpha,\beta,\gamma\in L^{\infty}((0,T);L^{\infty}(\Omega\times V))$, and consider $f^{0}_{j}\in L^{p}(\Omega\times V)$, for $p\geq 2$. Assume $Q^{(P)}\equiv 0$. Then, there is a unique solution $(f_{S},f_{I},f_{R})$ of the kinetic system (3), with $\varepsilon>0$, with $f_{j}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$, with initial condition given by $(f_{S}^{0},f_{I}^{0},f_{R}^{0})$.
Proof.
We assume without loss of generality that $\varepsilon=1$ and start with the existence part. Whenever we refer to a constant $C$ (with or without index), we consider that “there exists a positive number $C$ such that”. Its precise value is immaterial to the proof of the theorem, and the same parameter $C$ may refer to different constants along the proof.
Using the boundary condition $\hat{n}\cdot\nabla f_{j}|_{\partial\Omega}=0$, we conclude that
$0=\int_{\Omega}\nabla_{x}f_{j}^{p}\mathrm{d}x=p\int_{\Omega}f_{j}^{p-1}\nabla_{x}f_{j}\mathrm{d}x\ .$
We multiply Eq. (3) by $f_{j}^{p-1}$ and integrate over $\Omega\times V$:
$\frac{1}{p}\frac{\mathrm{d}\ }{\mathrm{d}t}\int_{\Omega\times V}f_{j}^{p}\mathrm{d}v\mathrm{d}x=\int_{\Omega\times V}f_{j}^{p-1}\left[\partial_{t}f_{j}+v\cdot\nabla_{x}f_{j}\right]\mathrm{d}v\mathrm{d}x=\int_{\Omega\times V}f_{j}^{p-1}\mathfrak{Q}_{j}[\mathbf{f}]\mathrm{d}v\mathrm{d}x$ (12)
We remember the Holder’s inequality:
$\int|gh|\leq\left(\int|g|^{q_{1}}\right)^{1/q_{1}}\left(\int|h|^{q_{2}}\right)^{1/q_{2}}\ ,\quad\frac{1}{q_{1}}+\frac{1}{q_{2}}=1\ ,\quad q_{1},q_{2}\geq 1\ .$
For $g=f_{j}\geq 0$ and $h=1$, with $q_{1}=p$, $q_{2}=\nicefrac{{p}}{{(p-1)}}$ we conclude that
$\int_{V}f_{j}\mathrm{d}v\leq|V|^{1-\frac{1}{p}}\left(\int_{V}f_{j}^{p}\mathrm{d}v\right)^{1/p}\ .$
For $g=f_{j}^{p-1}$, $h=1$, $q_{1}=\nicefrac{{p}}{{(p-1)}}$, $q_{2}=p$, we find
$\int_{V}f_{j}^{p-1}\mathrm{d}v\leq|V|^{\frac{1}{p}}\left(\int_{V}f_{j}^{p}\mathrm{d}v\right)^{1-\frac{1}{p}}\ .$
Multiplying the last two equations, we conclude
$\int_{\Omega\times V}f_{j}^{p-1}Q_{j}^{(R)}[\mathbf{f}]\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega}\lambda_{j}\left[\frac{1}{|V|}\int_{V\times V}f_{j}^{\prime}f_{j}^{p-1}\mathrm{d}v^{\prime}\mathrm{d}v-\int_{V}f_{j}^{p}\mathrm{d}v\right]\mathrm{d}x$
$=\int_{\Omega}\lambda_{j}\left[\frac{1}{|V|}\int_{V}f_{j}\mathrm{d}v\int_{V}f_{j}^{p-1}\mathrm{d}v-\int_{V}f_{j}^{p}\mathrm{d}v\right]\mathrm{d}x\leq 0\ .$ (13)
We proceed to prove the existence part of $f_{I}$. Note that
$\int_{\Omega\times V}Q^{(T)}_{I}f_{I}^{p-1}\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}\left[\frac{\beta f_{S}f_{I}^{p}}{f_{S}+f_{S}+f_{R}}-\gamma f_{I}^{p}\right]\mathrm{d}v\mathrm{d}x$
$\leq\|\beta\|_{L^{\infty}(\Omega\times V)}\|f_{I}\|_{L^{p}(\Omega\times V)}^{p}\ .$
Therefore
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{I}\|_{L^{p}(\Omega\times V)}\leq C\|f_{I}\|_{L^{p}(\Omega\times V)}\ .$
Using Gronwall’s inequality, we conclude that $f_{I}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$.
The next step is to show that
$\int_{\Omega\times V}Q^{(T)}_{R}f_{R}^{p-1}\mathrm{d}v\mathrm{d}x$ $=\int_{\Omega\times V}\left[\gamma f_{I}f_{R}^{p-1}-\alpha f_{R}^{p}\right]\mathrm{d}v\mathrm{d}x$
$\leq\|\gamma\|_{L^{\infty}(\Omega\times V)}\|f_{I}\|_{L^{p}(\Omega\times V)}\|f_{R}\|_{L^{p}(\Omega\times V)}^{p-1}\ .$
Using the estimate given by Eq. (12) and the existence result for $f_{I}$, we conclude
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|f_{R}\|_{L^{p}(\Omega\times V)}\leq C\ ,$
and, therefore $f_{R}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$.
We proceed similarly to prove that
$\int_{\Omega\times V}Q^{(T)}_{S}[\mathbf{f}]f_{S}^{p-1}\mathrm{d}v\mathrm{d}x\leq\|\alpha\|_{L^{\infty}(\Omega\times V)}\|f_{S}\|_{L^{p}(\Omega\times V)}^{p-1}\ ,$
and, therefore, $f_{S}\in L^{\infty}((0,T);L^{p}(\Omega\times V))$.
Now, we proceed to prove the uniqueness of the solution.
Let us define the norm
$\|\mathbf{f}-\mathbf{g}\|^{2}:=(f_{S}-g_{S})^{2}+(f_{I}-g_{I})^{2}+(f_{R}-g_{R})^{2}\ .$ (14)
From the linearity of $Q^{(R)}_{j}$ it is not difficult to prove that
$Q^{(R)}_{j}[\mathbf{f}]-Q^{(R)}_{j}[\mathbf{g}]=Q^{(R)}_{j}[\mathbf{f}-\mathbf{g}]\ .$
Therefore, from Eq. (13),
$\int_{\Omega\times V}\left(Q^{(R)}_{j}[\mathbf{f}]-Q^{(R)}_{j}[\mathbf{g}]\right)\left(f_{j}-g_{j}\right)\mathrm{d}v\mathrm{d}x=\int_{\Omega\times V}Q^{(R)}_{j}[\mathbf{f}-\mathbf{g}](f_{j}-g_{j})\leq 0\ .$ (15)
In the sequel, we will use the inequality $ab\leq\frac{1}{2}(a^{2}+b^{2})$.
We study the class-transition kernel:
$\left(Q^{(T)}_{S}[\mathbf{f}]-Q^{(T)}_{S}[\mathbf{g}]\right)\left(f_{S}-g_{S}\right)+\left(Q^{(T)}_{I}[\mathbf{f}]-Q^{(T)}_{I}[\mathbf{g}]\right)\left(f_{I}-g_{I}\right)$
$\qquad+\left(Q^{(T)}_{R}[\mathbf{f}]-Q^{(T)}_{R}[\mathbf{g}]\right)\left(f_{R}-g_{R}\right)$
$\quad=\left(-\beta\frac{f_{I}f_{S}}{f_{S}+f_{I}+f_{R}}+\alpha f_{R}+\beta\frac{g_{I}g_{S}}{g_{S}+g_{I}+g_{R}}-\alpha g_{R}\right)(f_{S}-g_{S})$
$\qquad+\left(\beta\frac{f_{I}f_{S}}{f_{S}+f_{I}+f_{R}}-\gamma f_{I}-\beta\frac{g_{I}g_{S}}{g_{S}+g_{I}+g_{R}}+\gamma g_{R}\right)(f_{I}-g_{I})$
$\qquad+\left(\gamma f_{I}-\alpha f_{R}-\gamma g_{I}+\alpha g_{R}\right)(f_{R}-g_{R})$
$\quad=\beta\left(\frac{f_{I}f_{S}}{f_{S}+f_{I}+f_{R}}-\frac{g_{I}g_{S}}{g_{S}+g_{I}+g_{R}}\right)\left((f_{I}-g_{I})-(f_{S}-g_{S})\right)$
$\qquad+\alpha(f_{R}-g_{R})(f_{S}-g_{S})-\gamma(f_{I}-g_{I})^{2}+\gamma(f_{I}-g_{I})(f_{R}-g_{R})$
$\qquad-\alpha(f_{R}-g_{R})^{2}\ .$
To simplify the notation, we define $A_{f}:=f_{S}+f_{I}+f_{R}$, and a similar notation for $A_{g}$. We prove that
$\left|\frac{f_{I}f_{S}}{A_{f}}-\frac{g_{I}g_{S}}{A_{g}}\right|$ $=\frac{1}{A_{f}A_{g}}\left|f_{S}g_{S}(f_{I}-g_{I})+f_{I}g_{I}(f_{S}-g_{S})+f_{S}g_{R}(f_{I}-g_{I})\right.$
$\quad\left.-f_{S}g_{I}(f_{R}-g_{R})+f_{R}g_{I}(f_{S}-g_{S})\right|$
$\leq C\left(|f_{S}-g_{S}|+|f_{I}-g_{I}|+|f_{R}-g_{R}|\right)\ .$
In the last passage, we used that
$\frac{f_{j}g_{j^{\prime}}}{A_{f}A_{g}}\leq 1\ .$
We conclude that
$\sum_{j=S,I,R}\left(Q^{(T)}_{j}[\mathbf{f}]-Q^{(T)}_{j}[\mathbf{g}]\right)\left(f_{j}-g_{j}\right)$ $\leq C_{1}(f_{S}-g_{S})^{2}+C_{2}(f_{I}-g_{I})^{2}+C_{3}(f_{R}-g_{R})^{2}$
$\leq C\|\mathbf{f}-\mathbf{g}\|^{2}\ .$
Considering the last equation and Eq, (15), we prove that
$\frac{\mathrm{d}\ }{\mathrm{d}t}\|\mathbf{f}-\mathbf{g}\|^{2}\leq 0\ ,$
and the uniqueness follows. ∎
Remark 4.
As discussed at the beginning of this section, we did not consider the preferred kernel in the proof of Thm. 2. As will be discussed in the conclusions, we believe that the existence and uniqueness theorem still holds with parameters $\eta,\xi>0$. If Thm. 2 could be extended such that $f_{I}\in L^{\infty}((0,T);H^{1}(\Omega\times V))$, then a bound would follow immediately.
We finish noting that the existence and uniqueness of solutions of the mesoscopic model (9)–(11), when $\Gamma\equiv 0$, can be proved in a similar way. However, we will not consider the case $\Gamma$ not identically zero, i.e., when S individuals actively try to escape from I individuals.
5 Examples
In this section, we consider three different examples. The first one, in Subsec. 5.1, is of a pedagogical nature and deals with a situation with constant parameters.
However, the reason to consider kinetic formulations of epidemic models is to introduce the velocity as a parameter, and this has interest only if we want to study the interaction between the pathogen dynamics and the mobility of individuals.
In the second example, in Subsec. 5.2, we consider a situation that unfortunately happens in several diseases when in certain localities, ranging from small villages to entire countries, large parts of the population behave as if the disease did not exist. Frequently, this is driven by misinformation campaigns. However, outside this region, susceptible individuals try to avoid infectious individuals, as is natural to expect.
Finally, Subsec. 5.3 considers that infectious individuals voluntarily confine themselves, represented by a decrease in the velocity of I individuals with respect to S and R individuals. In the previous discussion, the set of admissible velocities $V$ was the same for all classes of individuals, and considering a different set of admissible individuals introduces serious technical difficulties when dealing with the class-transition kernels $Q^{(T)}$. Therefore, we will consider this situation only in the one-dimensional case, in which formal calculations are significantly simpler.
5.1 The constant parameters case
Assume that $\alpha,\beta,\gamma,\eta,\xi$ are positive constants. In this case, $\kappa=vF(|v|)/\lambda$, and $\Theta=(\eta+\xi)vF(|v|)/\lambda$, and, therefore the coefficients of the mesoscopic equations are given by
$D=\frac{1}{\lambda}\int_{V}v^{2}F(|v|)\mathrm{d}v\ ,\qquad\Gamma=\frac{\eta+\xi}{\lambda}\int_{V}v^{2}F(|v|)\mathrm{d}v\ .$
The SIRS model will be, in this case, defined by parameters $\bar{\alpha}=\alpha$, $\bar{\beta}=\beta$, and $\bar{\gamma}=\gamma$, cf. Subsec. 3.3.
If we assume further that the space of admissible velocities is a circle or a sphere of radius $v_{0}$, for a movement in the two- or three-dimensional space, respectively, then $D=v_{0}^{2}/\lambda$ and $\Gamma=(\eta+\xi)v_{0}^{2}/\lambda$.
5.2 Different disease risk acceptances
Let us assume that there is a proper subset $\Omega_{0}\subset\Omega$ such that $\eta=\xi=\eta_{0}$ for $x\in\Omega_{0}$ and $\eta=\xi=\eta_{1}$ otherwise, with constants $\eta_{0}<\eta_{1}$. This means that in the region $\Omega_{0}$, S individuals are more risk-prone to the disease, with contrary behavior outside $\Omega_{0}$.
The interesting point is that the definition of $\kappa$ and $\Theta$, and, by extension, of $D$ and $\Gamma$ are local in $(t,x)$, and therefore the mesoscopic limits in each point $x\in\Omega$ will be obtained as if all $\eta$ and $\xi$ were constants in the entire domain with values equal to its values at $x$.
5.3 Confining infectious individuals
In this example, contrary to what was discussed so far, we consider that S, I, and R individuals have different velocity sets, allowing a possible confinement (voluntary or imposed) of I individuals. However, we will consider only the one-dimensional problem, as the dealing with class-transition kernels for different sets $V$ in the general multi-dimensional case is outside the scope of the present work.
We will assume three set {±$\{\pm v_{S}\}$, {±$\{\pm v_{I}\}$, {±$\{\pm v_{R}\}$, $v_{j}>0$, representing possible velocities for S, I, and R individuals, respectively. Density of individuals will be given by $f_{j}^{\pm}=f_{j}(t,x,\pm v_{j})$, with the superscript sign indicating if the corresponding individuals move to the right or to the left. Mesoscopic densities are given by $\rho_{j}=f_{j}^{-}+f_{j}^{+}$.
The equilibrium distribution is given by $F(v)=\frac{1}{2}$, and we proceed formally as in the Subsec. 3.2. In particular, the leading order is given by
$0=Q^{(R)}_{j}[\mathbf{f}]=\lambda_{j}\left(\frac{f_{j,0}^{+}+f_{j,0}^{-}}{2}-f_{j,0}^{\pm}\right)\ ,$
which implies that $f_{j,0}^{+}=f_{j,0}^{-}=\frac{\rho_{j}}{2}$. For the order 1 equation, we need to solve
$\lambda_{j}\left(\frac{\kappa^{+}_{j}-\kappa^{-}_{j}}{2}\right)=Q^{(R)}_{j}[\kappa_{j}]=-\frac{v_{j}}{2}\ ,$
on the assumption that $\kappa_{j}^{+}+\kappa_{j}^{-}=0$. We conclude that $\kappa_{j}^{\pm}=\pm\frac{v_{j}}{2\lambda_{j}}$, and, therefore $D_{j}=v_{j}^{2}/\lambda_{j}$. Finally,
$\Theta_{j}=\frac{(\eta+\xi)v_{j}}{2\lambda_{j}}\delta_{jS}\ ,$
where $\delta_{SS}=1$ and $\delta_{jS}=0$ otherwise. From this, we conclude that
$\Gamma_{j}=\frac{(\eta+\xi)v_{j}}{\lambda_{j}}\delta_{jS}\ .$
Mesoscopic equations are given by
$\partial_{t}\rho_{S}-\nabla_{x}\cdot\left(\frac{(\eta+\xi)v_{S}^{2}}{\lambda_{S}}(\nabla_{x}\rho_{I})\rho_{S}\right)-\nabla_{x}\cdot\left(\frac{v_{S}^{2}}{\lambda_{S}}\nabla_{x}\rho_{S}\right)=Q^{(T)}_{S}\left[\frac{\bm{\rho}}{2}\right]\ ,$
$\partial_{t}\rho_{j}-\nabla_{x}\left(\frac{v_{j}^{2}}{\lambda_{j}}\nabla_{x}\rho_{j}\right)=Q^{(T)}_{j}\left[\frac{\bm{\rho}}{2}\right]\ ,\qquad j=I,R\ ,$
where $\bm{\rho}=(\rho_{S},\rho_{I},\rho_{R})$.
We finish by noting that if $v_{I}\to 0$, then the diffusion coefficient of I individuals converges to 0. Therefore, in this limit case, we expect that if $\rho_{I}(0,x)=0$, then $\rho_{I}(t,x)=0$ for all $t>0$.
6 Conclusions and future work
The idea of this work is to provide a model compatible with the well-known SIRS model in mathematical epidemiology, but with more flexibility to include several effects that cannot be easily implemented in spatially-homogeneous compartmental models.
We considered that individuals move freely in a bounded domain and contagion occurs when susceptible and infectious individuals are walking side by side, i.e., with the same velocity and at the same point in space. We used recent ideas both in kinetic models and in game theory to provide a model that is compatible, at least at the formal level, with the SIRS system of ordinary differential equations.
An important point missing from the present work is to show the links between the several different descriptive scales. More precisely, one important result to be discussed in a future work is the convergence of the solutions of the kinetic model given by Eq. (3) to the solution of the mesoscopic drift-diffusion-advection model given by Eqs. (9), (10), and (11). Mathematically, this means to study in which space, for which parameters, and for which initial conditions can we guarantee $\lim_{\varepsilon\to 0}\int_{V}f_{j,\varepsilon}\mathrm{d}v=\rho_{j}$. In the same vein, we should study how small perturbations of homogeneous initial conditions behave in time, i.e., to discuss if the solution of the SIRS system of ordinary differential equations is a good approximation of the more realistic non-homogenous problem, both in the kinetic and in the mesoscopic (drift-diffusion-advection) level, when the problem is close to the homogenous one. In particular, several important parameters, such as the basic reproductive number $\mathcal{R}_{0}$, are defined in the homogeneous case. As there is no universally accepted formulation of a non-homogeneous compartmental model, it is unclear how robust these concepts are in providing reliable information when dealing with real data, which is naturally non-homogeneous. In [WZ12], a generalization of $\mathcal{R}_{0}$ was provided for non-homogeneous epidemic models at the PDE level. Yet, we are not aware of the definition of any similar concept at the kinetic level.
Another missing point of the present work is to consider in more detail the global existence and uniqueness of solutions when the preferred direction kernel is non-zero. One possible approach, that goes in the direction of [CMPS04, Model 2] is to replace $v\cdot\nabla_{x}f_{I}$ by $\frac{1}{\varepsilon}\left(f_{I}(x+\varepsilon v)-f_{I}(x)\right)$, with an appropriate high-order correction. In principle, this should not change the macroscopic limit, but it may be more manageable for rigorous proofs.
Assume that the preferred direction kernel is of one of the two forms discussed in the previous paragraph. From the modeling point of view, we expect $\eta,\xi>0$ to have a regularizing effect on the solution of the kinetic model. More precisely, assume that solution $\mathbf{f}$ exists in a given space for $\eta=\xi=0$. Considering that I individuals are generated by S individuals, and for $\eta,\xi>0$, S individuals tend to avoid high concentrations of I individuals, and considering that the absence of S individuals close to the high concentration of I will provide a limitation in the emergence of new I individuals, we expect that a solution would also exist in this new situation.
The situation would be completely different if we allowed $\eta$ and/or $\xi$ to be negative. For the problem we are interested in, this would not be realistic; however, in the more general game-theoretical situation discussed in Subsec. 3.3, that could be of interest, modeling, for example, the formation of aggregates starting from a small perturbation of a spatially-homogeneous solution. This idea is close to the original chemotaxis problem [KS70], which inspired the kinetic model discussed in [CMPS04], one of the main sources of the present work. The general kinetic model with game-theoretical class transition kernels approach could be of interest in the study of pattern formation mediated by games played between cells, cf. [Tom97, DCS+09] for applications of evolutionary game theory in cancer modeling, or in the evolution of cooperation between rational agents, cf. [AH81] for the initial studies in the evolution of cooperation and [NBM94] for its extension to spacial games.
Another limitation of the model is that individual behavior is purely mechanistic. However, typical human behavior considers at the same time local and global information, a level of detail that cannot be explored in ODE models. Furthermore, this information can be contradictory. An explicit example will help to understand that point. Consider a new class of individuals, the Vaccinated one, and assume that vaccination is voluntary. Human behavior depends on global information $\int_{\Omega\times V}f_{I}\mathrm{d}v\mathrm{d}x$ (i.e., the endemic situation in the entire population) and in the behavior of close friends and relatives (local information, i.e., $f_{V}$, the new class of type V individuals), and could also depend on the immediate perception, which is also a local information. These values can be contradictory, i.e., even if a large outbreak is happening, a cluster of unvaccinated individuals can exist, especially in isolated communities that were not yet reached by the outbreak. It is not clear that simple modifications of the kinetic or the PDE model discussed is the right model to capture that feature, but it is clear that an ODE model will not, as there is no meaningful difference between local and global in the last case. Human behavior and vaccination is studied from the mathematical point of view since the seminal work by [BE04]; see also [DRSC16, CDPdCS24].
In the near future, we hope to provide some answers for at least some of the questions above.
Acknowledgement
All the authors are funded by national funds through the FCT – Fundação para a Ciência e a Tecnologia, I.P., under the scope of the projects UIDB/ 00297/2020 (https://doi.org/10.54499/UIDB/00297/2020) and UIDP/00297/ 2020 (https:doi.org/10.54499/UIDP/00297/2020) (Center for Mathematics and Applications — NOVA Math) and the project Mathematical Modelling of Multi-scale Control Systems: applications to human diseases (CoSysM3) 2022.03091.PTDC (https://doi.org/10.54499/2022.03091.PTDC), supported by national funds (OE), through FCT/MCTES. All the authors thank Ana Jacinta Soares (Universidade do Minho, Portugal) for insightful comments on preliminary versions of the present work. FACCC also acknowledges Robert Pego (Carnegie Mellon University) for suggesting the application of the current theory in the 1D case, with varying velocity sets, resulting in the example discussed in Subsec. 5.3.
All the authors contributed equally to the development of the work’s ideas, computational codes, data analysis, discussions, and writing of the final version of the manuscript.